I got this number as my answer, but i m not sure if it is right. 0.01941556

Can you please solve the question again so i can compare my answer. thank u

A radioactive object emits particles according to a Poisson process at an average rate of 7.5 particles per second. We observe the object for a total of 5 seconds. Suppose it is known that at least 20 particles will be emitted during this interval, what is the probability that no more than 25 particles will be emitted during this interval?
Answer #1

for 5 seconds expected number of particles =7.5*5=37.5

hence P(no more than 25 particles given at least 20 particles)=P(X<=25|X>=20)

=P(20<=X<=25)/P(X>=20)=(P(X<=25)-P(X<=19))/(1-P(X<=19))=(0.02008-0.000664)/(1-0.000664)=0.0194285